OpenStudy (anonymous):

Name the property shown. 6 + 5 + 84 = 84 + 5 + 6 A. Associative property of addition B. Associative property of multiplication C. Commutative property of addition D. Commutative property of multiplication

OpenStudy (mathstudent55):

a + b = b + a commutative property (change of order) (a + b) + c = a + (b + c) associative property (change of grouping)

OpenStudy (anonymous):

A. Commutative Property of Addition

OpenStudy (anonymous):

Actually...it is not A, it is C

OpenStudy (anonymous):

Sorry that's what I meant, C.